OpenStudy (anonymous):

Solve equation x(x-2)(x+3)=0

jimthompson5910 (jim_thompson5910):

Hint: if AB = 0, then A = 0 or B = 0

jimthompson5910 (jim_thompson5910):

So this means that if x(x-2)(x+3)=0 then x=0, x-2=0, or x+3=0

jimthompson5910 (jim_thompson5910):

Solve each equation for x to find your three answers

OpenStudy (anonymous):

I got it to (x^2-2x)(x+3)=0

jimthompson5910 (jim_thompson5910):

no need to do that

OpenStudy (anonymous):

I just don't know how to distribute it

jimthompson5910 (jim_thompson5910):

it's valid of course, but you're going backwards

OpenStudy (anonymous):

that's the way my school wants me to do it though

jimthompson5910 (jim_thompson5910):

really? that's odd...

OpenStudy (anonymous):

oh wait

jimthompson5910 (jim_thompson5910):

do you see what I wrote above? that's the best way to do it

OpenStudy (anonymous):

yea, I understand

OpenStudy (anonymous):

how do you get the result then?

jimthompson5910 (jim_thompson5910):

alright great

jimthompson5910 (jim_thompson5910):

solve each equation for x

OpenStudy (anonymous):

I need a solution set

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

x-2=0 x = ???

jimthompson5910 (jim_thompson5910):

x+3=0 x = ???

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

good, and what else

OpenStudy (anonymous):

aren't I only suppose to get 2 answers for a solution set?

OpenStudy (anonymous):

wait wait wait

jimthompson5910 (jim_thompson5910):

no, in this case, you'll get 3

OpenStudy (anonymous):

if I get it to (x^2-2x)(x+3)=0, can I just divide the first one by 2 and get (x-2)(x+3)=0?

jimthompson5910 (jim_thompson5910):

no need to do that though

OpenStudy (anonymous):

so the answer is x=2 and x=-3?

jimthompson5910 (jim_thompson5910):

there's one more

OpenStudy (anonymous):

0?

jimthompson5910 (jim_thompson5910):

bingo

jimthompson5910 (jim_thompson5910):

x(x-2)(x+3)=0 x=0, x-2=0, or x+3=0 x=0, x=2, or x = -3

OpenStudy (anonymous):

woo!!

OpenStudy (anonymous):

thank you :3

jimthompson5910 (jim_thompson5910):

yw